Need some help soving this by graphing.
(Systems of linear equations)
3x+y=5
-2x+3y=4

Answer by mukhopadhyay(490)

3x+y=5..... (eq 1)
-2x+3y=4 ...(eq 2)
Multiply (eq 1) with 2 and (eq 2) with 3
6x+2y=10
-6x+9y=12
Adding both equations:
11y=22
=>y=2
3x+y=5
=>3x=5-y=5-2=3
=>x=1
Answer: x=1 and y=2